# ReHub: Linear Complexity Graph Transformers with Adaptive Hub-Spoke Reassignment

### v1

@@ A.1 Datasets
In Tables 5 and 6 we summaries the details of the datasets used for evaluation.
@@
PascalVOC-SP 11,355 5,443,545 479.40 5.65 30,777,444 2,710.48 10.74$\pm$0.51 27.62$\pm$2.13
COCO-SP 123,286 58,793,216 476.88 5.65 332,091,902 2,693.67 10.66$\pm$0.55 27.39$\pm$2.14
PCQM-Contact 529,434 15,955,687 30.14 2.03 32,341,644 61.09 4.63$\pm$0.63 9.86$\pm$1.79
Peptides-func 15,535 2,344,859 150.94 2.04 4,773,974 307.30 20.89$\pm$9.79 56.99$\pm$28.72
Peptides-struct 15,535 2,344,859 150.94 2.04 4,773,974 307.30 20.89$\pm$9.79 56.99$\pm$28.72
@@ A.2 Hyperparameters
Table 8: Model-specific hyperparameters for PCQM-Contact.
Hyperparameter # Layers Hidden Dim # Heads Hubs Ratio k
GCN 5 300 1 1 3
GCNII 5 100 2 1 3
GINE 5 100 1 1 3
GatedGCN 8 72 1 1 3
Table 9: Model-specific hyperparameters for Peptides-func.
Hyperparameter # Layers Hidden Dim # Heads Hubs Ratio k
GCN 5 155 1 0.5 3
GCNII 5 88 1 0.5 3
GINE 5 88 2 0.5 3
GatedGCN 5 88 1 0.5 3
Table 10: Model-specific hyperparameters for Peptides-struct.
Hyperparameter # Layers Hidden Dim # Heads Hubs Ratio k
GCN 5 155 1 0.5 3
GCNII 5 88 1 0.5 3
GINE 5 88 2 0.5 3
GatedGCN 5 88 1 0.5 3
@@
Table 12: Hyperparameters for OGBN-Arxiv and Coauthor Physics datasets used for evaluation.
@@ A.3 Implementation details
During preprocessing we use the METIS partitioning algorithm (Karypis & Kumar, 1998) to divide each graph to clusters according to the required number of hubs. In practice, we use the python wrapper PyMetis [^3] https://github.com/inducer/pymetis which allows us to map each spoke to a single hub.
@@ A.4 Additional ablations
$BC(P,Q)$ lies between 00 and $1$, where the higher the coefficient the more similar the distributions are. In our setup, $P$ is the distribution over spokes connection per hub; i.e. it is the number of spokes connected to each hub, divided by the overall number of connections from spokes to hubs (so that the result is indeed a probability distribution). We then set $Q$ to be the uniform distribution, i.e. $1/N_{h}$ for each hub. By doing so, we can see how close the actual distribution $P$ is to the uniform distribution, where uniform indicates the optimal balanced assignment of spokes to hub – where every hub has exactly the same number of spokes. For convenience, we define the Bhattacharyya Percentage as the Bhattacharyya Coefficient multiplied by $100$.

### v2

@@ A.1 Datasets
In Tables 6 and 7 we summaries the details of the datasets used for evaluation.
@@
[…]
@@ A.2 Hyperparameters
# Params - - - 590,051
Table 9: Model-specific hyperparameters for PCQM-Contact, and the number of model parameters.
Hyperparameter # Layers Hidden Dim # Heads Hubs Ratio k # Params
GCN 5 300 1 1 3 5,127,540
GCNII 5 100 2 1 3 587,240
GINE 5 100 1 1 3 638,240
GatedGCN 8 72 1 1 3 655,992
Table 10: Model-specific hyperparameters for Peptides-func, and the number of model parameters.
Hyperparameter # Layers Hidden Dim # Heads Hubs Ratio k # Params
GCN 5 155 1 0.5 3 1,372,563
GCNII 5 88 1 0.5 3 452,204
GINE 5 88 2 0.5 3 491,804
GatedGCN 5 88 1 0.5 3 611,044
Table 11: Model-specific hyperparameters for Peptides-struct, and the number of model parameters.
Hyperparameter # Layers Hidden Dim # Heads Hubs Ratio k # Params
GCN 5 155 1 0.5 3 1,376,382
GCNII 5 88 1 0.5 3 453,152
GINE 5 88 2 0.5 3 492,752
GatedGCN 5 88 1 0.5 3 611,992
@@
Table 13: Hyperparameters for OGBN-Arxiv and Coauthor Physics datasets used for evaluation, and the number of model parameters.
@@ A.3 Implementation details
During preprocessing we use the METIS partitioning algorithm (Karypis & Kumar, 1998) to divide each graph to clusters according to the required number of hubs. In practice, we use the python wrapper PyMetis [^4] https://github.com/inducer/pymetis which allows us to map each spoke to a single hub. According to the METIS paper (Karypis & Kumar, 1998), this clustering step runs in approximately $O(N+M)$ time and requires only $O(N)$ additional memory beyond the input graph, for a graph with $N$ nodes and $M$ edges. For graphs where $M\approx O(N)$, as in many practical cases, both the runtime and the additional memory simplify to $O(N)$. The clustering is implemented as a modular preprocessing stage and can, in principle, be replaced with any other clustering algorithm without affecting the overall ReHub framework.
@@
Balanced random assignment.
In Section 4.3 and Table 5, we present an ablation study of the spoke–hub assignment components. Specifically, we discuss a balanced random assignment strategy that balances the number of spokes assigned to each hub while keeping the number of hubs assigned to each spoke strictly equal to $k$. This strategy ensures per-spoke uniqueness and global balance, at the cost of introducing some structure into the randomness. Given $N_{h}$ hubs and $N_{s}$ spokes, we first sample a uniform random permutation $\pi$ of the hub indices $\{0,\dots,N_{h}{-}1\}$. We then choose a random stride $u\in\{1,\dots,N_{h}{-}1\}$ such that $\gcd(u,N_{h})=1$, ensuring a full cycle through the hubs. For each spoke $i\in\{0,\dots,N_{s}{-}1\}$, we assign it $k$ distinct hubs with indices
$E_{i,j}=\pi[(i\cdot u+j)\bmod N_{h}],\quad j=0,\dots,k{-}1.$
By construction, each spoke is assigned exactly $k$ distinct hubs, and each hub is assigned to approximately the same number of spokes. The partial randomness of $\pi$ and $u$ ensures diversity of assignments, while the cyclic structure maintains balanced and distinct connections per spoke.
Algorithm 2 Balanced Random Assignment of Hubs
0: Number of hubs $N_{h}$, number of spokes $N_{s}$, number of hubs per spoke $k$
 Sample a random permutation $\pi$ of $\{0,1,\dots,N_{h}{-}1\}$
 Sample a random stride $u\in\{1,\dots,N_{h}{-}1\}$ such that $\gcd(u,N_{h})=1$
for $i=0$ to $N_{s}-1$ do
  for $j=0$ to k1k-1 do
   $E_{i,j}=\pi[(i\cdot u+j)\bmod N_{h}\bigr{]}$
  end for
end for
return  Assignment matrix $\bm{E}$}
Why is it important for $u$ to be coprime with $N_{h}$? If $u$ and $N_{h}$ are coprime, then the sequence of values $(i\cdot u)\bmod N$ for $i=0,1,\dots,N-1$ forms a permutation of $\{0,1,\dots,N-1\}$. That is, each value in the sequence appears exactly once before the sequence repeats. To see this, suppose $(i\cdot u)\bmod N=(j\cdot u)\bmod N$ for some $0\leq i<j<N$. Then $(j-i)\cdot u$ is a multiple of $N$, but since $u$ and $N$ are coprime, jij-i must be a multiple of $N$. However, 0<ji<N0<j-i<N, a contradiction. Hence all $N$ values are distinct, and since they all lie in $\{0,\dots,N-1\}$, they must form a permutation of this set. This property ensures that all hubs are used evenly.
@@ A.4 Additional ablations
$BC(P,Q)$ lies between 0 and $1$, where the higher the coefficient the more similar the distributions are. In our setup, $P$ is the distribution over spokes connection per hub; i.e. it is the number of spokes connected to each hub, divided by the overall number of connections from spokes to hubs (so that the result is indeed a probability distribution). We then set $Q$ to be the uniform distribution, i.e. $1/N_{h}$ for each hub. By doing so, we can see how close the actual distribution $P$ is to the uniform distribution, where uniform indicates the optimal balanced assignment of spokes to hub – where every hub has exactly the same number of spokes. For convenience, we define the Bhattacharyya Percentage as the Bhattacharyya Coefficient multiplied by $100$.